if 25cm3 of a soltuion of NaOH are neutralised by 27.5cm3 of H2SO4 of a concentration of 0.250moldm-3 what is the concentration of the NaOH?


1
Expert's answer
2022-01-18T10:46:02-0500

Balanced chemical equation for the reaction:

2NaOH + H2SO4 products Na2SO4 + 2H2O

Provided data:

NaOH; volume - 25cm3 = 0.025dm3

H2SO4; volume - 27.5cm3 = 0.0275dm3

concentration - 0.250moldm-3

  • Applying the formula C=n/V;where n is moles,C is concentration and V is Volume,we then calculate the moles of 0.0275dm3 H2SO4 that reacted with NaOH.
  • n=CV

n (moles of H2SO4) = 0.250mol/dm3 H2SO4 × 0.0275dm3 H2SO4 (dm3 cancels out).

= 6.875 × 10-3 moles H2SO4

From the balanced equation,

Mole ratio of NaOH : H2SO4 is 2:1

Therefore,

6.875 × 10-3 moles H2SO4 × 2 moles NaOH/1 moles H2SO4

 

= 1.375 × 10-2 moles NaOH

  • Applying the formula C=n/V, we calculate the concentration of NaOH.

n=1.375 × 10-2 moles, V= 0.025 dm3

C= 1.375 × 10-2 moles/0.025 dm3 = 0.55 moldm-3

Therefore, concentration of NaOH is 0.55 moldm-3